A consumer buys 80 units of a good at price of ₹5 per unit. Suppose price elasticity of demand is (-) 2 . At what price will he buy 64 units?


1
Expert's answer
2021-10-01T09:17:47-0400

"E_d=\frac{\Delta Q}{\Delta P}\times \frac{P}{Q}"

"\Delta Q= 64 -80=-16"


"-2=\frac{-16}{\Delta P}\times \frac{5}{80}"

"\Delta P(-2)= -1"

"\Delta P= 0.5"


New price = original price+ change in price.

"New Price= 5 +0.5= 5.5"

"= 5.5" per unit